# Generating Chain-of-Thoughts with a Direct Pairwise-Comparison Approach to Searching for the Most Promising Intermediate Thought

### v1

@@ 4 Experiments
We test our proposed algorithm in three real-world tasks: question answering, as well as mathematical reasoning tasks, namely, the Game of 24 and Sudoku Puzzles. The LLM employed in experiments is GPT-3.5-turbo-1106  [^2] https://platform.openai.com/docs/models.
@@ 4.1 Experiment Setup
    Comp-SToT: We replace the score-based evaluation in the original SToT algorithm with our proposed direct pairwise comparison approach and refer to this variant algorithm as Comp-SToT.
@@
Intermediate Thoughts Generation. In general, different tasks should have different thought generators. Exploiting problem properties is essential to effectively design the intermediate thoughts. An intermediate thought should be “small” and “simple” enough for LLMs to generate promising and diverse thoughts. We follow the setting of Yao et al. (2023) to generate the intermediate thoughts. For example, we generate the thoughts as a few words, as in QA; as a line of equations, as in the Game of 24; or as an intermediate solution in the Sudoku puzzle.
We defer the implementation details for prompts and the cost comparison of our proposed approaches and other contenders to Appendix A and Appendix C.
@@ 4.2 Question Answering
Task Setup. We first test the performance of our proposed algorithm on the Question Answering (QA) task using the AQuA dataset (Ling et al., 2017), which comprises 254 arithmetic reasoning tasks aimed at assessing logical abilities through various mathematical computation problems. Each question in this dataset is accompanied by five multiple-choice options, labeled from A to E. We follow the experimental protocol as it is in the work of Wang et al. (2022). The accuracy of the responses is gauged by comparing the generated answers with the standard solutions.
@@
Method Accuracy Direct 24.8% CoT 42.3% SC-CoT 58.4% SToT 57.1% SC-SToT 57.6% Comp-SToT 59.0% Back-SToT 58.0% C-ToT (Stand.) 61.4% C-ToT (Duel.) 63.0% Table 1: Average Accuracy Comparison on AQuA.
Figure 3: Prediction Details of SToT and Comp-SToT on AQuA.
Comparison Results. We report the comparison results of our proposed approaches with other contenders in Table 1. All CoT approaches outperform the Direct query method, showing the importance of designing effective CoTs to guide LLMs from simplicity to complexity. We can also observe that the SC-SToT method outperforms the original SToT method, where the algorithm scores the intermediate thoughts multiple times to alleviate the noise. However, this mechanism will significantly increase the token cost. We leave the detailed discussion of token cost to Appnedix C.
Both proposed variants of SToT methods achieve higher average accuracy than the original SToT method. In Comp-SToT, the point-wise scoring mechanism is replaced by a direct pairwise comparison, while in Back-SToT, our backtracking search algorithm replaces the original search algorithm, taking into account all previous intermediate thoughts. These results demonstrate the effectiveness of these two mechanisms, such that the proposed C-ToT (Stand.) and C-ToT (Deul.) approaches outperform all contenders. Moreover, C-ToT (Deul.) achieves superior performance by further modeling noise in the comparison.
We delve deeper to explore the benefits of the direct pairwise comparison mechanism to test whether it can better find the most promising intermediate thoughts. Note that we do not have access to the underlying value or order of the intermediate thoughts in each round. Therefore, we use the final prediction error as a proxy, since the depth of the tree structure in the QA datasets is shallow, limited to 1 to 3 levels. We quantify the number of QA problems correctly/incorrectly predicted by ToT and Comp-SToT and report it in Figure 3. Our observation shows that Comp-SToT predicts more correctly when S-ToT predicts incorrectly, showing the superiority of the pairwise comparison mechanism over the pointwise scoring mechanism. This validates the rationale of direct pairwise comparison.
@@ 4.3 Game of 24
C-ToT Setup. In this task, we restore unselected intermediate thoughts in previous layers and apply pruning when the current node is inferior to previously unselected thoughts. We set the maximum depth of tree of thoughts to 6. The computation terminates either when an answer containing the number 24 is derived, or when the maximum layer limit is reached.
@@
We also optimize the pruning process that takes place before the comparison stage. Any newly generated intermediate thought with a single number unequal to 24 is eliminated. The rest are then filtered by comparison, selecting a number of new thoughts equal to or less than the maximum number of selected thoughts. Thoughts that are not selected are added to the “remain list”. If one of the selected thoughts contains the final answer, it is added to the “answer list” and the interactive process is stopped. We leave the implementation details of the thought generator and the comparison prompt to Appendix A.2.
Method Accuracy Direct 8.0% CoT 4.3% SC-CoT 8.0% SToT 34.3% SC-SToT 40.0% Comp-SToT 36.3% Back-SToT 39.0% C-ToT (Stand.) 40.0% C-ToT (Duel.) 41.0% Table 2: Average Accuracy Comparison on Game of 24.
Figure 4: Accuracy of different methods against their token costs.
Comparison Results. We report the overall comparison results in Table 2. We observed trends similar to those reported in the QA task. We find that both the SToT and CToT approaches significantly outperform the CoT-based methods, indicating the need to interact with the LLM to generate a more powerful chain of thoughts to handle complex reasoning tasks. The SC-SToT method improves the performance of SToT, while the Comp-SToT can achieve a similar improvement with direct pairwise comparison, indicating the superiority of the comparison mechanism. Therefore, the combination of the comparison mechanism and the specific design for comparison noise allows the C-ToT (Duel.) to achieve the highest average accuracy.
We also report the average accuracy of different methods against their token cost in Figure 4. We can observe that the token cost per question of the Comp-SToT algorithm is the same as that of the SToT algorithm, but it achieves a better performance. In practice, we can reduce the token cost by using a counter for each thought to track its comparison frequency. If the comparison count of a thought exceeds a threshold, we can remove it from the tree. We also leave the detailed discussion of token cost to Appnedix C.
@@ 4.4 Sudoku Puzzle
Task Setup. In this task, we use the Sudoku dataset  [^3] https://github.com/jieyilong/tree-of-thought-puzzle-solver/data/benchmarks/sudoku, which contains 10 Sudoku puzzles each of 3x3, 4x4, and 5x5 dimensions. Each puzzle is partially filled with some numbers, and the task is to complete the entire Sudoku grid without changing the given numbers. The correctness of the solutions was determined by whether a complete and correct Sudoku grid is generated.
@@
Method Acc. $3\times 3$ Acc. $4\times 4$ Acc. $5\times 5$ Direct 56.7% 37.7% 16.7% CoT 73.3% 36.7% 23.3% SC-CoT 76.7% 50.0% 16.7% SToT 86.7% 46.7% 46.7% SC-SToT 96.7% 53.3% 50.0% Comp-SToT 100.0% 46.7% 50.0% Back-SToT 100.0% 60.0% 56.7% C-ToT (Stand.) 100.0% 63.3% 60.0% C-ToT (Duel.) 100.0% 63.3% 63.3% Table 3: Average Accuracy Comparison on the Sudoku Puzzle.
Comparison Results. We report the overall comparison results in Table 3. In all three tasks, our proposed approaches, C-ToT (Stand.) and C-ToT (Duel.), consistently achieve the highest average accuracy, demonstrating their superior ability to handle complex reasoning tasks. While the SToT method generally outperforms the Direct method and the CoT method, it does not always outperform the SC-ToT method, as seen in the $4\times 4$ Sudoku task. This phenomenon may be due to the potential noise in pointwise scoring methods, which could mislead the subsequent generation of intermediate thoughts. The SC-ToT method introduces thought ensembles, which naturally mitigate the noise in the LLM feedback. Our proposed methods consistently outperform other contenders, suggesting that direct pairwise comparison effectively mitigates noise in LLM feedback, identifies the most promising intermediate thoughts, and improves the generated chain compared to SToT-based methods. We can observe that the C-ToT (Stand.) algorithm achieves the same performance as the C-ToT (Duel.) algorithm, which indicates that a single direct pairwise comparison could already provide reliable feedback in the Sudoku puzzle tasks.
@@ 5 Conclusion
This paper investigates a widespread but understudied problem of noisy feedback from LLMs in CoT generation tasks. Motivated by Vapnik’s principle, we argue that for LLMs, the simultaneous comparison of two thoughts provides a more robust evaluation compared to individual value evaluations, and thus we propose a direct pairwise-Comparison ToT approach C-ToT to find the most promising intermediate thought. The proposed method directly searches for the most promising intermediate thought by pairwise comparison, and incorporates previous thoughts into the comparison to allow for rethinking. To further alleviate the noise in the comparison, we investigate two noise models, propose two variants of the C-ToT algorithm, and analyze the theoretical properties. Experiments on three real-world mathematical and reasoning tasks show the effectiveness of our proposed algorithm and verify the rationale of direct pairwise comparison.
A Implementation Details
@@ B Proofs
Lemma 2 (Lemma 2 in Falahatgar et al. (2017)).
@@
Proof  To make this paper self-contained, we provide the proofs in Falahatgar et al. (2017) here. First, we prove that the probability of the direct pairwise comparison process providing a wrong winner is less than $\delta$. Let $\widehat{p}_{i}^{r}$ and $\widehat{c}^{r}$ denote $\widehat{p}_{i}$ and $\widehat{c}$ respectively after $r$ number of comparisons. The output of the pairwise comparison will not be $i$ only if $\widehat{p}_{i}^{r}<\frac{1}{2}+\varepsilon-\widehat{c}^{r}$ for any $r<m=\frac{1}{2\varepsilon^{2}}\log\frac{2}{\delta}$ or if $\widehat{p_{i}}<\frac{1}{2}$ for $r=m$.
@@
$Pr(\widehat{p}_{i}^{r}<\frac{1}{2}+\varepsilon-\widehat{c}^{r})\leq e^{-2r(\widehat{c}^{r})^{2}}=e^{-\log\frac{4r^{2}}{\delta}}=\frac{\delta}{4r^{2}}.$
@@
$Pr(\exists r\text{ s.t. }\widehat{p}_{i}^{r}\leq\frac{1}{2}+\varepsilon-\widehat{c}^{r})\leq\frac{\delta}{2}.$
@@
As each of the $|Z|/2$ pairs is compared at most $\frac{1}{2\varepsilon^{2}}\log\frac{2}{\delta}$ times, the total comparisons is $\leq\frac{|Z|}{4\varepsilon^{2}}\log\frac{2}{\delta}$. Let $k^{*}$ be the comparison winner and $z^{*}$ be the maximum in $Z$. Let $a$ be the element paired with $z^{*}$. There are two cases: $\widetilde{p}(z^{*},a)\geq\varepsilon$ and $\widetilde{p}(z^{*},a)<\varepsilon$.
@@
where (a) follows from $r(a)\leq r(z^{*})$, (b) and (c) follow from the definitions of $z^{*}$ and $k^{*}$ respectively. From strong stochastic tranisitivity on $a$, $k^{*}$ and $z^{*}$, $\widetilde{p}(z^{*},k^{*})\leq\gamma\widetilde{p}(z^{*},a)\leq\gamma\varepsilon$.
@@
Proof [of Lemma 1] To make this paper self-contained, we also provide the proofs in Falahatgar et al. (2017) here. We first show that with probability $\geq 1-\delta$, the output of Knockout is an $\varepsilon$-maximum. Let $\varepsilon_{i}=c\varepsilon/2^{i/3}$ and $\delta_{i}=\delta/2^{i}$. Note that $\varepsilon_{i}$ and $\delta_{i}$ are bias and confidence values used in round $i$. Let $b_{i}$ be a maximum element in the set $Z$ before round $i$. Then by Lemma 2, with probability $\geq 1-\delta_{i}$,
@@
Proof [of Proposition 1] In each round of comparison, according to Lemma 1, we have the probability of $1-\delta$ to output the $\varepsilon$-maximum thoughts. For the intermediate thoughts in the $\tau$-th layer in the tree of depth $T$, the probability of them being the $\varepsilon$-maximum thoughts but not being selected is $1-\delta^{(T-\tau+1)}$.
@@
Therefore, the probability of missing the $\varepsilon$-maximum promising thoughts in the $\tau$-th layer is $1-\delta^{(T-\tau+1)}$ with $\mathcal{O}\left(\frac{\gamma^{2}n\sum_{i=1}^{T}|Z^{i}|}{\varepsilon^{2}}\log\frac{1}{\delta}\right)$ comparisons.
C Cost and Efficiency

### v2

@@ 4 Experiments
We test our proposed algorithm in three real-world tasks: question answering (QA), as well as mathematical reasoning tasks, namely, the Game of 24 and Sudoku Puzzles. The LLM employed in experiments is GPT-3.5-turbo-1106.
@@ 4.1 Experiment Setup
    Comp-SToT: We replace the score-based evaluation in the original SToT algorithm with our proposed pairwise comparison approach and refer to this variant algorithm as Comp-SToT.
@@
In addition, we include three state-of-the-art algorithms into comparison: PoT (Chen et al., 2023), Self-Refine (Madaan et al., 2023) and GoT (Besta et al., 2023). Detailed experimental results can be found in Appendix C.1.
@@
Intermediate Thoughts Generation. In general, different tasks should have different thought generators. Exploiting problem properties is essential to effectively design the intermediate thoughts. We follow the setting of (Yao et al., 2023) to generate the intermediate thoughts. For example, we generate the thoughts as a few words, as in QA; as a line of equations, as in the Game of 24; or as an intermediate solution in the Sudoku puzzle. We defer the implementation for prompts and the cost comparison of our approaches and other contenders to Appendix A and Appendix C.2.
@@ 4.2 Question Answering
Task Setup. We first test the performance of our proposed algorithm on the question answering tasks using the AQuA dataset (Ling et al., 2017), which comprises 254 arithmetic reasoning tasks aimed at assessing logical abilities through various mathematical computation problems. Each question in this dataset is accompanied by five multiple-choice options, labeled from A to E. We follow the experimental protocol as it is in the work of Wang et al. (2022). The accuracy of the responses is gauged by comparing the generated answers with the standard solutions. Results on other QA datasets can be found in Appendix C.1.
@@
Method Accuracy
Direct 24.8%
CoT 42.3%
SC-CoT 58.4%
SToT 57.1%
SC-SToT 57.6%
Comp-SToT 59.0%
Back-SToT 58.0%
C-ToT (Stand.) 61.4%
C-ToT (Duel.) 63.0%
Table 1: Average accuracy on AQuA.
Figure 3: Predictions of SToT and Comp-SToT on AQuA.
Comparison Results. We report the comparison results of our proposed approaches with other contenders in Table 1. All CoT approaches outperform the Direct query method, showing the importance of designing effective CoTs to guide LLMs from simplicity to complexity. We can also observe that the SC-SToT method outperforms the original SToT method, where the algorithm scores the intermediate thoughts multiple times to alleviate the noise. However, this mechanism will significantly increase the token cost. We leave the detailed discussion of token cost to Appnedix C.2.
Both proposed variants of SToT methods achieve higher average accuracy than the original SToT method. In Comp-SToT, the point-wise scoring mechanism is replaced by a pairwise comparison, while in Back-SToT, our backtracking search algorithm replaces the original search algorithm, taking into account all previous intermediate thoughts. These results demonstrate the effectiveness of these two mechanisms, such that the proposed C-ToT (Stand.) and C-ToT (Duel.) approaches outperform all contenders. Moreover, C-ToT (Duel.) achieves superior performance by further modeling noise in the comparison.
We delve deeper to explore the benefits of the pairwise comparison mechanism to test whether it can better find the most promising intermediate thoughts. Note that we do not have access to the underlying value or order of the intermediate thoughts in each round. Therefore, we use the final prediction error as a proxy, since the depth of the tree structure in the QA datasets is shallow, limited to 1 to 3 levels. We quantify the number of QA problems correctly/incorrectly predicted by ToT and Comp-SToT and report it in Figure 3. Our observation shows that Comp-SToT predicts more correctly when S-ToT predicts incorrectly, showing the superiority of the pairwise comparison mechanism over the pointwise scoring mechanism. This validates the rationale of pairwise comparison mechanism.
@@ 4.3 Game of 24
C-ToT Setup. In this task, we restore unselected intermediate thoughts in previous layers and apply pruning when the current node is inferior to previously unselected ones. We set the maximum depth of tree to 6. The computation terminates either when an answer containing the number 24 is derived, or when the maximum layer limit is reached.
@@
We also optimize the pruning process that takes place before the comparison stage and apply it to all contenders to save tokens. Newly generated intermediate thoughts with a single number unequal to 24 is eliminated. The rest are then filtered by comparison, selecting a number of new thoughts equal to or less than the maximum number of selected thoughts. Thoughts that are not selected are added to the “remain list”. If one of the selected thoughts contains the final answer, it is added to the “answer list” and the interactive process is stopped. We leave the implementation details of the thought generator and the comparison prompt to Appendix A.2.
Method Accuracy
Direct 8.0%
CoT 4.3%
SC-CoT 8.0%
SToT 34.3%
SC-SToT 40.0%
Comp-SToT 36.3%
Back-SToT 39.0%
C-ToT (Stand.) 40.0%
C-ToT (Duel.) 41.0%
Table 2: Average accuracy on Game of 24.
Figure 4: Accuracy and token costs of differenta methods.
Comparison Results. We report the overall comparison results in Table 2. We observed trends similar to those reported in the QA task. We find that both the SToT and CToT approaches significantly outperform the CoT-based methods, indicating the need to interact with the LLM to generate a more powerful chain of thoughts to handle complex reasoning tasks. The SC-SToT method improves the performance of SToT, while the Comp-SToT can achieve a similar improvement with pairwise comparison, indicating the superiority of the comparison mechanism. Therefore, the combination of the comparison mechanism and the specific design for comparison noise allows the C-ToT (Duel.) to achieve the highest average accuracy.
We also report the average accuracy of different methods against their token cost in Figure 4. We can observe that the token cost per question of the Comp-SToT algorithm is the same as that of the SToT algorithm, but it achieves a better performance. In practice, we can reduce the token cost by using a counter for each thought to track its comparison frequency. If the comparison count of a thought exceeds a threshold, we can remove it from the tree. We also leave the detailed discussion of token cost to Appnedix C.2.
@@ 4.4 Sudoku Puzzle
Task Setup. We use the Sudoku Dataset (Long, 2023), containing 10 Sudoku puzzles each of 3x3, 4x4, and 5x5 dimensions. Each puzzle is partially filled with numbers, and the task is to complete the entire Sudoku grid without changing the given numbers. The correctness of the solutions was determined by whether a complete and correct Sudoku grid is generated.
@@
Method Acc. $3\times 3$ Acc. $4\times 4$ Acc. $5\times 5$
Direct 56.7% 37.7% 16.7%
CoT 73.3% 36.7% 23.3%
SC-CoT 76.7% 50.0% 16.7%
SToT 86.7% 46.7% 46.7%
SC-SToT 96.7% 53.3% 50.0%
Comp-SToT 100.0% 46.7% 50.0%
Back-SToT 100.0% 60.0% 56.7%
C-ToT (Stand.) 100.0% 63.3% 60.0%
C-ToT (Duel.) 100.0% 63.3% 63.3%
Table 3: Average accuracy on Sudoku Puzzles.
Comparison Results. We report the comparison results in Table 3. In all three tasks, our proposed approaches, C-ToT (Stand.) and C-ToT (Duel.), consistently achieve the highest average accuracy, demonstrating their superior ability to handle complex reasoning tasks. While the SToT method generally outperforms the Direct method and the CoT method, it does not always outperform the SC-ToT method, as seen in the $4\times 4$ Sudoku task. This phenomenon may be due to the potential noise in pointwise scoring methods, which could mislead the subsequent generation of intermediate thoughts. The SC-ToT method introduces thought ensembles, which naturally mitigate the noise in the LLM feedback. Our proposed methods consistently outperform other contenders, suggesting that pairwise comparison mechanism effectively mitigates noise in LLM feedback, identifies the most promising intermediate thoughts, and improves the generated chain compared to SToT-based methods. We can observe that the C-ToT (Stand.) algorithm achieves the same performance as the C-ToT (Duel.) algorithm, which indicates that a single pairwise comparison could already provide reliable feedback in the Sudoku puzzle tasks.
@@ 5 Conclusion
This paper investigates a widespread but understudied problem of noisy feedback from LLMs in CoT generation tasks. Motivated by Vapnik’s principle, we argue that for LLMs, the simultaneous comparison of two thoughts provides a more robust evaluation compared to individual value evaluations, and thus we propose a pairwise-comparison ToT approach C-ToT, approaching to searching for the most promising intermediate thought. The proposed method directly selects the most promising intermediate thought by pairwise comparison, and incorporates previous thoughts into the comparison to allow for rethinking. To further alleviate the noise in the comparison, we propose two variants of the C-ToT algorithm, and analyze the theoretical properties. Experiments on three real-world mathematical and reasoning tasks show the effectiveness of our proposed algorithm and verify the rationale of the pairwise comparison.
Acknowledgments
HY was supported by Cisco Faculty Research Award. MS was supported by JST CREST Grant Number JPMJCR18A2.
Impact Statement
This research investigates a general problem of CoT generation with any LLM, where we take into account the noise in the feedback of the LLM. Therefore, when using LLMs for complex mathematical or logical reasoning problems, the user could benefit from our study from the aspect of generating a more effective CoT. The consequences of system failure and bias in the data are not applicable.
Appendix A Implementation Details
@@ Appendix B Proofs
Lemma 2 (Lemma 2 in (Falahatgar et al., 2017)).
@@
Proof of Lemma 2.
To make this paper self-contained, we provide the proofs in (Falahatgar et al., 2017) here. First, we prove that the probability of the direct pairwise comparison process providing a wrong winner is less than $\delta$. Let $\widehat{p}_{i}^{r}$ and $\widehat{c}^{r}$ denote $\widehat{p}_{i}$ and $\widehat{c}$ respectively after $r$ number of comparisons. The output of the pairwise comparison will not be $i$ only if $\widehat{p}_{i}^{r}<\frac{1}{2}+\varepsilon-\widehat{c}^{r}$ for any $r<m=\frac{1}{2\varepsilon^{2}}\log\frac{2}{\delta}$ or if $\widehat{p_{i}}<\frac{1}{2}$ for $r=m$.
@@
$\Pr(\widehat{p}_{i}^{r}<\frac{1}{2}+\varepsilon-\widehat{c}^{r})\leq e^{-2r(\widehat{c}^{r})^{2}}=e^{-\log\frac{4r^{2}}{\delta}}=\frac{\delta}{4r^{2}}.$
@@
$\Pr(\exists r\text{ s.t. }\widehat{p}_{i}^{r}\leq\frac{1}{2}+\varepsilon-\widehat{c}^{r})\leq\frac{\delta}{2}.$
@@
As each of the $|Z|/2$ pairs is compared at most $\frac{1}{2\varepsilon^{2}}\log\frac{2}{\delta}$ times, the total comparisons is less than $\frac{|Z|}{4\varepsilon^{2}}\log\frac{2}{\delta}$. Let $k^{*}$ be the comparison winner and $z^{*}$ be the maximum in $Z$. Let $a$ be the element paired with $z^{*}$. There are two cases: $\widetilde{p}(z^{*},a)\geq\varepsilon$ and $\widetilde{p}(z^{*},a)<\varepsilon$.
@@
where (a) follows from $r(a)\leq r(z^{*})$, (b) and (c) follow from the definitions of $z^{*}$ and $k^{*}$ respectively. From strong stochastic tranisitivity on $a$, $k^{*}$ and $z^{*}$, $\widetilde{p}(z^{*},k^{*})\leq\gamma\widetilde{p}(z^{*},a)\leq\gamma\varepsilon$. ∎
@@
Proof of Lemma 1.
To make this paper self-contained, we also provide the proofs in (Falahatgar et al., 2017) here. We first show that with probability $\geq 1-\delta$, the output of Knockout is an $\varepsilon$-maximum. Let $\varepsilon_{i}=c\varepsilon/2^{i/3}$ and $\delta_{i}=\delta/2^{i}$. Note that $\varepsilon_{i}$ and $\delta_{i}$ are bias and confidence values used in round $i$. Let $b_{i}$ be a maximum element in the set $Z$ before round $i$. Then by Lemma 2, with probability $\geq 1-\delta_{i}$,
@@
Proof of Proposition 1.
In each round of comparison, according to Lemma 1, we have the probability of $1-\delta$ to output the $\varepsilon$-maximum thoughts. Suppose that the thoughts in the shallower layers are more promising than those in the deeper ones, which is often the case in step-by-step reasoning tasks. For example, $r([\mathbf{z}^{1}_{j}])\geq r([\mathbf{z}^{1}_{i},\mathbf{z}^{2}])$ for $j\in[K]$, $i\not\in[K]$, and $\mathbf{z}^{2}\in Z^{2}$. When generating the intermediate thoughts in the $\tau$-th layer, the probability that the $\varepsilon$-maximum thoughts are not selected is $1-\delta^{\tau}$.
@@
Therefore, the probability of missing the $\varepsilon$-maximum promising thoughts in the $\tau$-th layer is $1-\delta^{\tau}$ with at most $\mathcal{O}\left(\frac{\gamma^{2}\sum_{i=1}^{T}|Z^{i}|}{\varepsilon^{2}}\log\frac{1}{\delta}\right)$ comparisons required for generating the tree. ∎
Appendix C Additional Experimental Results
C.1 Experimental Results Summary
First, we present a summary of the experimental results.
We introduce three additional QA datasets—Gsm8k (Cobbe et al., 2021), Coin Flip (OOD) (Wei et al., 2022), and BBH (Srivastava et al., 2023)—along with two more state-of-the-art algorithms for comparison, with implementation details provided below.
Contenders. PoT (Chen et al., 2023): PoT requires in-context samples to guide LLMs generating Python code step-by-step, and the number of samples is a hyperparameter. Since SToT and C-ToT do not always require such samples, we choose appropriate number of samples in PoT to maintain experimental consistency with other methods. We use one in-context sample for all datasets expecting Game of 24, which uses 3 samples.
Self-Refine (Madaan et al., 2023): The parameter of in Self-Refine is the number of iterations of Self-Refine. For each task, we set the number of iterations to 3 so that the number of tokens used is approximately the same. We use the same template as in the original paper and set the number of in-context examples to the same as in the PoT.
We can observe in Table 4 that the proposed C-ToT (Stand.) and C-ToT (Duel.) algorithms achieve the best performance in almost all tasks. The Self-Refine also achieves promising results in QA tasks, while it performs relatively poorly compared to the SToT and C-ToT algorithms in complex tasks that require step-by-step interaction and reasoning, such as Game of 24.
Method / Data Gsm8k Coin flip (OOD) BBH AQuA Game of 24 Sudoku
CoT 68.8 $\pm$ 2.5 56.3 $\pm$ 2.1 67.7 $\pm$ 2.6 42.3 $\pm$ 2.5 4.3 $\pm$ 3.2 44.4 $\pm$ 2.3
SToT 59.3 $\pm$ 2.4 62.1 $\pm$ 2.9 69.6 $\pm$ 2.0 57.1 $\pm$ 2.1 34.3 $\pm$ 3.9 60.0 $\pm$ 3.8
PoT 62.2 $\pm$ 3.3 71.1 $\pm$ 2.8 66.7 $\pm$ 2.3 47.5 $\pm$ 4.5 27.2 $\pm$ 3.7 43.1 $\pm$ 3.9
Self-Refine 67.6 $\pm$ 3.2 64.8 $\pm$ 1.6 69.2 $\pm$ 3.7 56.2 $\pm$ 4.1 16.3 $\pm$ 3.9 52.3 $\pm$ 3.0
C-ToT (Stand.) 71.3 $\pm$ 3.3 66.2 $\pm$ 2.8 75.7 $\pm$ 1.9 61.4 $\pm$ 2.9 40.0 $\pm$ 4.6 74.4 $\pm$ 2.1
C-ToT (Duel.) 73.0 $\pm$ 2.7 70.4 $\pm$ 2.1 80.0 $\pm$ 2.2 63.0 $\pm$ 2.6 41.0 $\pm$ 3.0 75.4 $\pm$ 3.3
Table 4: Performance comparisons on benchmark datasets. On each dataset, 5 test runs were conducted and the average accuracy as well as standard deviation are presented, and the best one is emphasized in bold.
Our method is generally not comparable to GoT (Besta et al., 2023), but the idea of comparison can improve GoT. The GoT divides a complex task into several subtasks, generate thoughts for each subtask and merge them, while the ToT-style methods aim at step-by-step thinking, where the new generated thought is based on previous ones, so it is hard to split or merge. Thus, they are better suited for different tasks. For example, for sorting, we can divide it into subtasks; while for the Game of 24, the answer should be generated based on previous thoughts, so C-ToT is more proper.
We can use the idea of comparison to improve the solving of subtasks, so we can subsequently improve the GoT. The following is a preliminary study of the sorting task in the GoT paper, remaining all its setups, and comparing the performance with scoring mechanism by LLM and the pairwise comparison mechanism by LLM in subtasks. Accuracy and token usage are reported as in Table 5. 5 test runs were conducted and the average accuracy and the number of tokens (completion tokens / prompt tokens) are reported.
Data / Method S-GoT C-GoT
32 elements 90.4%; 1232/12726 90.5%; 1217/14969
64 elements 86.1%; 2592/31118 86.6%; 2356/35313
Table 5: Performance comparisons with GoT.
C.2 Cost and Efficiency
In Table 67 and 8, we report the token cost and average accuracy comparison of our proposed approaches with other contenders. 5 test runs were conducted and the average accuracy and the number of tokens (completion tokens / prompt tokens) are reported.